Required information

[The following information applies to the questions displayed below.]

Caro Manufacturing has two production departments, Machining and Assembly, and two service departments, Maintenance and Cafeteria. Direct costs for each department and the proportion of service costs used by the various departments for the month of August follow:

Proportion of Services Used by
Department Direct Costs Maintenance Cafeteria Machining Assembly
Machining $ 104,000
Assembly 74,400
Maintenance 46,000 0.2 0.5 0.3
Cafeteria 37,000 0.6 0.2 0.2

Required:

Compute the allocation of service department costs to producing departments using the direct method. (Do not round intermediate calculations.)

machining

assembly
Answer #1

Maintenance department costs will be divided between Machining and Assembly in the ratio of 5:3 and Cafeteria department costs will be divided between Machining and Assembly in the ratio of 1:1

Allocation of service department costs to producing departments

Maintenance department Cafeteria department Total
Machining 46,000 x 5/8 = $28,750 37,000 x 1/2 = $18,500 $47,250
Assembly 46,000 x 3/8 = $17,250 37,000 x 1/2 = $18,500 $35,750
Total $46,000 $37,000 $83,000

Please ask if you have any query related to the question. Thank you